You are given the following information about a closed economy with no government:
Consumption = 115 + 0.6Y
Investment = 550
Use the above information to answer the questions that follow:
Q.4.1 Calculate the value of autonomous spending.
Q.4.2 Calculate the value of the multiplier.
Q.4.3 Calculate the equilibrium level of income.
Q.4.4 Is the equilibrium level of income also the full employment level of income? Explain (3)
your answer.
Q.4.5 Identify any three non‐income determinants of consumption
1
Expert's answer
2020-10-02T06:29:24-0400


autonomous Spending, that part of spending, which is independent of income level

"115+550

= \$ 665"

4.2 Calculate the value of the multiplier.

"Multiplier m = 1\/(1-MPC)"

"= 1\/.4= 2.5"



4.3 Calculate the equilibrium level of income.

"Y=C+I"

"Y=115+0.6Y+550"

"Y = 115+550+0.6YY=115+550+0.6Y"

"Y=665+0.6Y"

"Y\u22120.6Y=665"

"0.4Y=665"

"Y = \\frac{665}{0.4}Y= \n0.4\n665\n\u200b"

"=1663"

4.4 Is the equilibrium level of income also the full employment level of income? Explain your answer.

yes, eqm income level is full employment level

4.5 Identify any three non‐income determinants of consumption

1) Expectations about future price level

2) real interest rates

3) price level & wealth